OpenStudy (anonymous):

Prove the identity: (1-sinx)(secx+tanx)=1 over secx P.s. Happy new years everyone!!! :)

OpenStudy (anonymous):

happy new years

OpenStudy (turingtest):

convert secx to 1/cosx convert 1/secx to cos x convert tanx to sinx/cosx then just multiply. You will wind up with 1-sin^2x in the numerator. Rewrite that as cos^2 and it's done.

OpenStudy (anonymous):

sec is h/a tan is o/a sin is o/h

OpenStudy (turingtest):

thank you sunsetlove...

OpenStudy (anonymous):

sec is h/a tan is o/a sin is o/h ?

OpenStudy (turingtest):

forget that, just listen to my post. can you do that? what do you get?

OpenStudy (anonymous):

so, (1-sinx)(1 over cosx + sinx over cosx) and if I multiply..

OpenStudy (turingtest):

yes... see what you get

OpenStudy (anonymous):

ohh \[1-\sin ^{2} \theta \over \cos \theta \] !??

OpenStudy (turingtest):

good and what is 1-sin^2x=??? (very important identity)

OpenStudy (anonymous):

\[\cos ^{2}x\]

OpenStudy (turingtest):

so the fraction becomes...?

OpenStudy (anonymous):

gotcha! :) wooooo la la. Thanks ;)

OpenStudy (anonymous):

cosx which is same as the right side :)

OpenStudy (turingtest):

Nice job Hannah!

OpenStudy (anonymous):

I have to say you are very helpful! :)

OpenStudy (turingtest):

I try, thanks for listening :)

OpenStudy (anonymous):

btw, did memorize all of these identities?

OpenStudy (anonymous):

did u *